OpenStudy (anonymous):

Travis is riding the Ferris wheel at the amusement park. His height can be modeled by the equation H(t) = 22 cos pi over 13t + 28, where H represents the height of the person above the ground in feet at t seconds.

OpenStudy (anonymous):

1. Part 1: How far above the ground is Travis before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Travis begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Travis' height above the ground reaches a minimum?

OpenStudy (anonymous):

@aryandecoolest can you help?

OpenStudy (anonymous):

oh!ok!! before ride begins. t=0 so \[22 \cos (\pi/13)*0 + 28\] ------>22+28=50

OpenStudy (anonymous):

2. Time period=\[2\pi/B\] \[B=\pi/18\] so time period for one revolution is \[2π/(π/13)=26 \sec\]

OpenStudy (anonymous):

there above.... B=pi/13, my bad.

OpenStudy (anonymous):

I notice that but I knew what you meant, but where does the 2pi come from, from 2π/(π/13)=26sec

OpenStudy (anonymous):

That's a relation.. The equation is in from of \[y=Acos(Bx-C)+D\] \[amplitude=22\] \[ B=π/13\]

OpenStudy (anonymous):

Oh ok, I see

OpenStudy (anonymous):

in part 3, do I combine part 1 and part 2 answers?

OpenStudy (anonymous):

you know when the cos function has minimum value @morningskye123

OpenStudy (anonymous):

is that a question? if it is im not to sure, trig- is weird and difficult. Is the mini going to be in (bx - c)?

OpenStudy (anonymous):

it will be minimum at t=1/2----> t=26/2 =13 sec, so from above equation we can also co relate value for vertical shift=28. ACtually vertical shift is the measure of how far the graph has shifted vertically, either up or down, from its initial position... so anyways, at t=13 \[H(13)=22\cos(π/13*13)+vertical shift \] \[H(13)=22*(-1)+ 28=6\] so it gives 6ft.

OpenStudy (anonymous):

Omg, thank you and that makes a lot of sense. So, thank you again

OpenStudy (anonymous):

lol!! np) anytime. :)

OpenStudy (anonymous):

Here's a metal :D